# Diversity of low-mass planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v2

@@ 4.1 Earth-sized planets
For Earth-sized planets, our results show that the dominant atmosphere at temperate conditions that emerges after planet formation is CO2-rich with a total surface pressure around 100 bar. Such an atmosphere forming upon cooling on the early Earth from a magma ocean state was first hypothesised by Holland (1984), computed by Zahnle et al. (2007) and demonstrated by Sossi et al. (2020) for a magma ocean initially equilibrated at its estimated O2 close to the IW buffer. CO2-rich atmospheres are compatible with that of Venus at the present day, both in composition and total pressure. The composition is also similar to that of modern Mars, hinting that this mode of atmosphere formation may have been commonplace on rocky planets in our solar system. That the Earth does not currently possess a 97:3 CO2–N2 atmosphere points to the operation of geochemical cycles that drew down carbon from the atmosphere through weathering by the Urey reaction to form carbonates (Sleep & Zahnle, 2001; Zahnle et al., 2007) that are subsequently recycled into the interior, thus evolving its atmosphere away from CO2-rich over geological timescales. This is because carbon is abundant, CO2 is strongly favoured at 280 K, and water condensation buffers the emergence of hydrogen-dominated atmospheres even when total C/H is low (as with atmospheres derived from a partially molten starting state, Appendix B).
@@
Condensates at planetary surfaces, such as water, graphite, -sulfur, and ammonium chloride, are of primary interest for the establishment of geochemical cycles (Hirschmann, 2018; Dasgupta et al., 2024) and a habitable environment (Benner et al., 2020; Mrnjavac et al., 2023; Rimmer & Shorttle, 2024). Figures 9 and B5 show atmospheres at 280 K above surfaces of an Earth-like planet together with stable condensates with which they are in equilibrium, derived from a fully molten and partially molten mantle, respectively. Each panel mandates that a particular condensate is stable with a minimum mass (shown on the x-axis), where the upper row reveals the number of models associated with a different dominant (VMR50%) species as a function of the number of Earth-ocean equivalents (EOs) condensed. The lower row shows the variety of atmospheric species for dominant classes of atmospheres that support condensates, where either CO2 or CH4 is chosen as the dominant species (Table 3). The conditions for water oceans and graphite to be stable are comparable, with two classes of atmospheres emerging as determined by a volume mixing ratio greater than 50%: CO2-rich and CH4-rich, with the former representing 83% of all atmospheres, whereas the latter only around 5% (e.g., Figure 9). For both classes, the atmospheres have median total surface pressures of a few tens of bar, and due to buffering imposed by C(cr) and H2O(l) condensates, the partial pressures of other C- and H-bearing species are relatively constant, regardless of the condensate reservoir mass.
@@
Small amounts of water can be stable beneath atmospheres dominated by SO2, N2, H2, and ClH, although these are all less prevalent than CO2 and CH4 (Figure 9a). Conditions that permit the condensation of solid ammonium chloride, ClH4N(cr) mirror those for water and graphite, reflecting the greater availability of NH3 and HCl in reduced atmospheres relative to their oxidised counterparts (Figure 2). Ammonium chloride most commonly condenses in small mass fractions (0.0025 EOs) beneath CO2-rich atmospheres, but the highest ClH4N(cr) mass fractions (0.006 up to 0.012 EOs) occur, albeit rarely (1% of all models), in equilibrium with H2. Conversely, -sulfur stability occurs almost exclusively beneath CO2-rich atmospheres, in which the partial pressure of CO2 is on average higher than in equilibrium with water and graphite (compare orange line in Figure 9e and g). The CO2-rich atmospheres in equilibrium with -sulfur also frequently have several tens of bar of SO2 and only 10-11 bar of H2S. Note that SO2(g) is almost never the major gas phase above -sulfur, as most of the S budget is hosted in the condensed phase. In total, water, graphite, -sulfur, and ammonium chloride coexist in around 40% of the models, albeit at potentially low abundance. Nevertheless, our results show that many early atmospheres are conducive to the formation of condensates on temperate planetary surfaces (here 280 K), should these atmospheres be able to cool isochemically and reach chemical equilibrium with stable condensates.
@@
Figure 10 summarises the atmospheric abundances (ppm) above planetary surfaces that simultaneously support water, graphite, -sulfur, and ammonium chloride. The objective is to identify characteristic atmospheric signatures that may be used to identify planetary surfaces with a higher propensity for habitability (Krissansen-Totton et al., 2018; Schwieterman et al., 2019; Konrad et al., 2024; Tokadjian et al., 2024; Borges et al., 2024; Young et al., 2024a). To satisfy joint condensate stability, reduced species, such as H2, NH3 and CH4 are low in abundance—almost always below the ppm level—because otherwise -sulfur does not form. This comes from the fact that these species are present in the atmosphere only at very low O2 at the high temperature initial conditions, at which S is highly soluble. Therefore, S is never present in sufficient amounts in these atmospheres to be able to condense upon cooling to a low temperature. Conversely, oxidised initial conditions that produce significant SO2 lead to final, elevated O2s of these atmospheres that prevent the condensation of graphite via CO2(g) = C(cr) + O2(g), even at high CO2. Consequently, any detection of SO2 in an atmosphere is incompatible with the presence of graphite. On the other hand, SO2 is highly unlikely to be the dominant atmospheric gas in the presence of H2O(l) or -S(cr). Species whose fugacities are less redox sensitive cannot readily be used to discriminate between stable versus unstable condensates (e.g., H2O and N2). Figure B6 illustrates the same analysis but excludes -sulfur and ammonium chloride and hence only considers the joint stability of water and graphite. In this scenario, reduced species can be accommodated at higher abundances (e.g. H2, NH3), and notably CH4-rich atmospheres are compatible with both water oceans and graphite at planetary surfaces.
@@
Our analysis provides a geological baseline for the expected states of planetary atmospheres after they emerge from the magma ocean stage, post-formation but prior to modification by geochemical cycling, atmospheric (photo)chemistry and escape, and later, possibly life. It should be noted that these computations assume chemical equilibrium is reached at 280 K in the absence of chemical modification of the atmosphere following the magma ocean stage. In practice, the timescales over which the atmospheres cool may be sufficiently short so as to permit chemical kinetics and photochemical processes (Arney et al., 2016; Zahnle et al., 2020; Drant et al., 2025), in addition to continued volcanic outgassing (Ortenzi et al., 2020; Guimond et al., 2021; Liggins et al., 2022) and/or atmospheric escape/erosion (e.g., Krissansen-Totton et al., 2024) to invalidate our dual assumptions of chemical equilibrium at a constant bulk composition. Hence, these results should be regarded as an end-member case, or to provide the initial atmospheric composition to inform the subsequent time evolution of cooling. We anticipate that our analysis can inform a prior expectation in modelling studies and statistical analyses, in which observations can be used to quantify whether a planetary atmosphere has evolved away from its earliest outgassed state.
@@
The establishment of geological cycling may require an early reservoir of water at the surface or in the upper mantle (e.g., Tikoo & Elkins-Tanton, 2017) to alter the rheology and partial melting of rocks (Korenaga, 2020). In this regard, surfaces that emerge post-formation with condensates, notably liquid water, may already be primed to cycle volatiles with the interior and thereby modify the size and composition of the atmosphere. Conversely, for surfaces that emerge without condensates, there could be a higher propensity to remain locked in an inactive state since geochemical cycling cannot initiate, at least not immediately. Table 3 reveals that most (91%) Earth-like planets that emerge from the magma ocean stage have the prerequisite conditions for surficial water, with at least 56% of planets also accommodating graphite. Since condensates buffer the atmospheric speciation (due to the equivalence of chemical potentials, an intensive variable), finding atmospheres that are compatible with stable condensates will not directly allow the size of those condensate reservoirs to be determined.
@@ 4.2 Sub-Neptunes
Our analysis permits new insights on two key aspects of the nature of sub-Neptunes; 1) the influence of atmospheres on their mass-radius relationships and demographics (e.g., Bean et al., 2021; Luque & Pallé, 2022) and 2) the identity and abundance of gas species in equilibrium with magma oceans, providing grounds for comparison with spectroscopic observational constraints (e.g., Madhusudhan et al., 2023; Benneke et al., 2024).
4.2.1 Non-ideality and the ‘fugacity crisis’
@@
Regarding point 1, Kite et al. (2019) proposed that a ‘fugacity crisis’, in which the enhanced dissolution of H2 at high surface pressures (1–8 GPa) hinders atmospheric growth, can explain the drop-off in exoplanet abundance above 3 R⊕ observed in exoplanet demographics studies (i.e., the ‘radius cliff’) (Fulton & Petigura, 2018); though see also Schlichting & Young (2022) for a different interpretation. Our work goes beyond the H2 non-ideality case considered in Kite et al. (2019), by incorporating additional volatile species (e.g., H2O, CH4, CO and CO2) along with their dissolution into the interior and real gas EOS. Despite the relatively large mass of envelopes (2–3% of the planetary mass) required to explain the mass-radius relationships of sub-Neptune exoplanets in the absence of interior dissolution, these quantities are modest, on a molar basis, with respect to the proportion of condensed mass in the planet. As such, in the end-member case of a solar gas interacting with a rocky sub-Neptune interior, the equilibrium H/O and C/O ratios will be modified from their solar values (2000 and 0.55 by moles, respectively, Asplund et al., 2009) via the mutual solubility of O in the envelope and C and H in the mantle. Since the H/O and C/O ratios of nominally rocky material are 1 (H/O and C/O ratios in the bulk silicate Earth are 0.005 and 3.6 10-4 by moles, respectively, Palme & O’Neill, 2014), chemical equilibrium will result in a decrease in H/O and C/O in the envelope relative to solar values (see also Young et al., 2023).
@@
Here, because only volatile element mass balance is considered (that is, there is no notion of the chemical composition of the condensed phase other than for C, H, O and N), this process is simulated by considering a higher O2 (IW=-3) in our sub-Neptune models than that imposed by the solar nebula gas (IW=-6.5, Grossman et al., 2008). Should O2 at the magma ocean-envelope interface fall below IW=-3.5, then H2 predominates in the gas phase (envelope), even for C/H ratios 100 solar, though CH4 remains a minor species (Figure G1). As noted by Kite et al. (2019); Schlichting & Young (2022), however, the experiments underpinning the solubility law for H2 dissolution in silicate liquids were performed over a limited temperature (1673–1773 K) and pressure (up to 3 GPa) range (Hirschmann et al., 2012); the temperature range is notably much lower than the 3000 K assumed for equilibrium at the interface. Therefore, how the equilibrium constant of the reaction H2(g) = H2(l) depends on temperature, and pressures above 3 GPa, is unconstrained, yet is essential to assess the ‘fugacity crisis’ hypothesis.
4.2.2 Role of non-ideality in inferring atmospheric chemistry
@@
A key feature of spectroscopic determinations of the speciation of gases in the atmospheres of temperature sub-Neptunes with equilibrium temperatures in the range 250–400 K, such as K2-18b ( = 255 K) and TOI-270d ( = 354 K) is the ubiquity of CO2(g) and CH4(g) (Madhusudhan et al., 2023; Benneke et al., 2024). Mixing ratios of these species are of the order 1–2% each, and their ratio, is roughly unity (Madhusudhan et al., 2023; Benneke et al., 2024), with evidence for H2O(g) at similar levels for TOI-270d but unresolved (below 10-3) in K2-18b. In all cases, the background gas is inferred to be H2. Although these transmission spectroscopy measurements probe the upper atmosphere (near 10-2–10-4 bar, Benneke et al., 2024; Madhusudhan et al., 2025) and the speciation cannot therefore be directly compared with that predicted in our models, they provide first-order constraints on the chemical degrees of freedom permitted in sub-Neptune atmospheres.
@@
The lower C/O ratios observed than predicted by our models leads to another discrepancy; the absence of CO2(g) in our simulations. As highlighted for terrestrial planetary atmospheres (Section 3.1), however, any CO(g) that is stable at high temperatures (here, 3000 K) is partitioned between CO2(g), C(cr) and/or CH4(g) at low temperatures (300–400 K, the temperatures in the photosphere recovered from transmission spectra) depending on the prevailing gas chemistry and total pressure. This implies substantial amounts of CO2(g) could form in the upper atmosphere upon cooling. Therefore, future work should concentrate on how gas-condensed phase equilibria act to modify the speciation of sub-Neptune atmospheres throughout the atmospheric column.
A simplified version of this exercise was carried out by Yang & Hu (2024), in which speciation at the pressures of observation in the atmospheres of K2-18b and TOI-270d were extrapolated to 200 bar. In order to match the CH4/CO2 proportions in the upper atmosphere, these authors inferred H2/H2O ratios at 200 bar to be roughly unity for K2-18b and 0.5 for TOI-270d. Assuming these ratios are representative of those at any magma ocean/atmosphere interface, then this would correspond to IW in the range 0 to 1. These O2s are similar to, or more oxidised than, those inferred for small, differentiated rocky bodies in the solar system (Wadhwa, 2008). Such high oxygen fugacities would seem at odds with the substantial quantities (1–3% by mass) of solar-like gas inferred to have been accreted by sub-Neptunes. Indeed, we predict C/H ratios that are too high (2.5) at these O2s, and H2 would no longer be a predominant constituent of the atmosphere (Figure G1). Therefore, preliminary indications suggest that either 1) bulk planetary C/H ratios should be relatively low (10-1 by mass), either by processes that fractionate C from other atmosphere-forming elements during accretion relative to solar or its preferential sequestration into sub-Neptune cores and/or 2) the lower atmosphere has a distinct chemical composition with higher C/H and C/O than does the upper atmosphere sensed by transmission measurements (e.g., by inefficient mixing or condensation of graphite in the atmospheric column), in order to maintain simultaneously low C/H (0.1–0.3) and C/O (0.02–0.5) ratios observed in temperate sub-Neptunes.

### v3

@@ 4.1 Earth-sized planets
For Earth-sized planets, our results show that the dominant atmosphere at temperate conditions that emerges after planet formation is CO2-rich with a total surface pressure around 100 bar. Such an atmosphere forming upon cooling on the early Earth from a magma ocean state was first hypothesized by Holland (1984), computed by Zahnle et al. (2007), and demonstrated by Sossi et al. (2020) for a magma ocean initially equilibrated at its estimated O2 close to the IW buffer. CO2-rich atmospheres are compatible with that of Venus at the present day, both in composition and total pressure. The composition is also similar to that of modern Mars, hinting that this mode of atmosphere formation may have been commonplace on rocky planets in our solar system. That the Earth does not currently possess a 97:3 CO2–N2 atmosphere points to the operation of geochemical cycles that drew down carbon from the atmosphere through weathering by the Urey reaction to form carbonates (Sleep & Zahnle, 2001; Zahnle et al., 2007) that are subsequently recycled into the interior, thus evolving its atmosphere away from CO2-rich over geological timescales. This is because carbon is abundant, CO2 is strongly favored at 280 K, and water condensation buffers the emergence of hydrogen-dominated atmospheres even when total C/H is low (as with atmospheres derived from a partially molten starting state, Appendix B).
@@
Condensates at planetary surfaces, such as water, graphite, -sulfur, and ammonium chloride, are of primary interest for the establishment of geochemical cycles (Hirschmann, 2018; Dasgupta et al., 2024) and a habitable environment (Benner et al., 2020; Mrnjavac et al., 2023; Rimmer & Shorttle, 2024). Figures 9 and B5 show atmospheres at 280 K above surfaces of an Earth-like planet together with stable condensates with which they are in equilibrium, derived from a fully molten and partially molten mantle, respectively. Each panel mandates that a particular condensate is stable with a minimum mass (shown on the -axis), where the upper row reveals the number of models associated with a different dominant (VMR 50%) species as a function of the number of Earth-ocean equivalents by mass (EO) condensed. The lower row shows the variety of atmospheric species for dominant classes of atmospheres that support condensates, where either CO2 or CH4 is chosen as the dominant species (Table 3). The conditions for water oceans and graphite to be stable are comparable, with two classes of atmospheres emerging as determined by a VMR greater than 50%: CO2-rich and CH4-rich, with the former representing 83% of all atmospheres, whereas the latter only around 5% (e.g., Figure 9). For both classes, the atmospheres have median total surface pressures of a few tens of bar, and due to buffering imposed by C(cr) and H2O(l) condensates, the partial pressures of other C- and H-bearing species are relatively constant, regardless of the condensate reservoir mass.
@@
Small amounts of water can be stable beneath atmospheres dominated by SO2, N2, H2, and ClH, although these are all less prevalent than CO2 and CH4 (Figure 9a). Conditions that permit the condensation of solid ammonium chloride, ClH4N(cr) mirror those for water and graphite, reflecting the greater availability of NH3 and HCl in reduced atmospheres relative to their oxidized counterparts (Figure 2). Ammonium chloride most commonly condenses in small mass fractions (0.0025 EO) beneath CO2-rich atmospheres, but the highest ClH4N(cr) mass fractions (0.006 up to 0.012 EO) occur, albeit rarely (1% of all models), in equilibrium with H2. Conversely, -sulfur stability occurs almost exclusively beneath CO2-rich atmospheres, in which the partial pressure of CO2 is on average higher than in equilibrium with water and graphite (compare the orange lines in Figure 9e and g). The CO2-rich atmospheres in equilibrium with -sulfur also frequently have several tens of bar of SO2 and only 10-11 bar of H2S. Note that SO2(g) is almost never the major gas phase above -sulfur because most of the S budget is hosted in the condensed phase. In total, water, graphite, -sulfur, and ammonium chloride coexist in around 40% of the models, albeit at potentially low abundance. Nevertheless, our results show that many early atmospheres are conducive to the formation of condensates on temperate planetary surfaces (here 280 K), should these atmospheres be able to cool isochemically and reach chemical equilibrium with stable condensates.
@@
Figure 10 summarizes the atmospheric abundances (ppm) above planetary surfaces that simultaneously support water, graphite, -sulfur, and ammonium chloride. The objective is to identify characteristic atmospheric signatures that may be used to identify planetary surfaces with a higher propensity for habitability (Krissansen-Totton et al., 2018; Schwieterman et al., 2019; Konrad et al., 2024; Tokadjian et al., 2024; Borges et al., 2024; Young et al., 2024a). To satisfy joint condensate stability, reduced species, such as H2, NH3 and CH4 are low in abundance—almost always below the ppm level—because otherwise -sulfur does not form. This comes from the fact that these species are present in the atmosphere only at very low O2 at the high-temperature initial conditions, at which S is highly soluble. Therefore, S is never present in sufficient amounts in these atmospheres to be able to condense upon cooling to a low temperature. Conversely, oxidized initial conditions that produce significant SO2 lead to final, elevated O2s of these atmospheres that prevent the condensation of graphite via CO2(g) = C(cr) + O2(g), even at high CO2. Consequently, any detection of SO2 in an atmosphere is incompatible with the presence of graphite. On the other hand, SO2 is highly unlikely to be the dominant atmospheric gas in the presence of H2O(l) or -S(cr). Species whose fugacities are less redox sensitive cannot readily be used to discriminate between stable versus unstable condensates (e.g., H2O and N2). Figure B6 illustrates the same analysis but excludes -sulfur and ammonium chloride, and hence only considers the joint stability of water and graphite. In this scenario, reduced species can be accommodated at higher abundances (e.g. H2, NH3), and notably CH4-rich atmospheres are compatible with both water oceans and graphite at planetary surfaces.
@@
Our analysis provides a geological baseline for the expected states of planetary atmospheres after they emerge from the magma ocean stage, postformation but prior to modification by geochemical cycling, atmospheric (photo)chemistry and escape, and later, possibly life. It should be noted that these computations assume chemical equilibrium is reached at 280 K in the absence of chemical modification of the atmosphere following the magma ocean stage. In practice, the timescales over which the atmospheres cool may be sufficiently short so as to permit chemical kinetics and photochemical processes (Arney et al., 2016; Zahnle et al., 2020; Drant et al., 2025), in addition to continued volcanic outgassing (Ortenzi et al., 2020; Guimond et al., 2021; Liggins et al., 2022) and/or atmospheric escape/erosion (e.g., Krissansen-Totton et al., 2024) to invalidate our dual assumptions of chemical equilibrium at a constant bulk composition. Hence, these results should be regarded as an end-member case, or to provide the initial atmospheric composition to inform the subsequent time evolution of cooling. We anticipate that our analysis can inform a prior expectation in modeling studies and statistical analyses, in which observations can be used to quantify whether a planetary atmosphere has evolved away from its earliest outgassed state.
@@
The establishment of geological cycling may require an early reservoir of water at the surface or in the upper mantle (e.g., Tikoo & Elkins-Tanton, 2017) to alter the rheology and partial melting of rocks (Korenaga, 2020). In this regard, surfaces that emerge postformation with condensates, notably liquid water, may already be primed to cycle volatiles with the interior and thereby modify the size and composition of the atmosphere. Conversely, for surfaces that emerge without condensates, there could be a higher propensity to remain locked in an inactive state since geochemical cycling cannot initiate, at least not immediately. Table 3 reveals that most (91%) Earth-like planets that emerge from the magma ocean stage have the prerequisite conditions for surficial water, with at least 56% of planets also accommodating graphite. Since condensates buffer the atmospheric speciation (due to the equivalence of chemical potentials, an intensive variable), finding atmospheres that are compatible with stable condensates will not directly allow the size of those condensate reservoirs to be determined.
@@ 4.2 Sub-Neptunes
Our analysis permits new insights on two key aspects of the nature of sub-Neptunes: (1) the influence of atmospheres on their mass-radius relationships and demographics (e.g., Bean et al., 2021; Luque & Pallé, 2022) and (2) the identity and abundance of gas species in equilibrium with magma oceans, providing grounds for comparison with spectroscopic observational constraints (e.g., Madhusudhan et al., 2023; Benneke et al., 2024).
4.2.1 Nonideality and the ”Fugacity Crisis”
@@
Regarding point 1, Kite et al. (2019) proposed that a ”fugacity crisis,” in which the enhanced dissolution of H2 at high surface pressures (1–8 GPa) hinders atmospheric growth, can explain the drop-off in exoplanet abundance above 3 observed in exoplanet demographics studies (i.e., the ”radius cliff”) (Fulton & Petigura, 2018); though see also Schlichting & Young (2022) for a different interpretation. Our work goes beyond the H2 nonideality case considered in Kite et al. (2019), by incorporating additional volatile species (e.g., H2O, CH4, CO, and CO2) along with their dissolution into the interior and real gas EOS. Despite the relatively large mass of envelopes (2%–3% of the planetary mass) required to explain the mass-radius relationships of sub-Neptune exoplanets in the absence of interior dissolution, these quantities are modest, on a molar basis, with respect to the proportion of condensed mass in the planet. As such, in the end-member case of a solar gas interacting with a rocky sub-Neptune interior, the equilibrium H/O and C/O ratios will be modified from their solar values (2000 and 0.55 by moles, respectively, Asplund et al., 2009) via the mutual solubility of O in the envelope and C and H in the mantle. Since the H/O and C/O ratios of nominally rocky material are 1 (H/O and C/O ratios in the BSE are 0.005 and 3.6 10-4 by moles, respectively, Palme & O’Neill, 2014), chemical equilibrium will result in a decrease in H/O and C/O in the envelope relative to solar values (see also Young et al., 2023).
@@
Here, because only volatile element mass balance is considered (i.e., there is no notion of the chemical composition of the condensed phase other than for C, H, O, and N), this process is simulated by considering a higher O2 (IW = -3) in our sub-Neptune models than that imposed by the solar nebula gas (IW = -6.5, Grossman et al., 2008). Should O2 at the magma ocean–envelope interface fall below IW = -3.5, then H2 predominates in the gas phase (envelope), even for C/H ratios 100 solar, though CH4 remains a minor species (Figure G1). As noted by Kite et al. (2019); Schlichting & Young (2022), however, the experiments underpinning the solubility law for H2 dissolution in silicate liquids were performed over a limited temperature (1673–1773 K) and pressure (up to 3 GPa) range (Hirschmann et al., 2012); the temperature range is notably much lower than the 3000 K assumed for equilibrium at the interface. Therefore, how the equilibrium constant of the reaction H2(g) = H2(l) depends on temperature, and pressures above 3 GPa, is unconstrained, yet is essential to assess the ”fugacity crisis” hypothesis.
4.2.2 Role of Nonideality in Inferring Atmospheric Chemistry
@@
A key feature of spectroscopic determinations of the speciation of gases in the atmospheres of temperature sub-Neptunes with equilibrium temperatures in the range 250–400 K, such as K2-18b ( = 255 K) and TOI-270d ( = 354 K), is the ubiquity of CO2(g) and CH4(g) (Madhusudhan et al., 2023; Benneke et al., 2024). Mixing ratios of these species are of the order 1–2% each, and their ratio, is roughly unity (Madhusudhan et al., 2023; Benneke et al., 2024), with evidence for H2O(g) at similar levels for TOI-270d but unresolved (below 10-3) in K2-18b. In all cases, the background gas is inferred to be H2. Although these transmission spectroscopy measurements probe the upper atmosphere (near 10-2–10-4 bar, Benneke et al., 2024; Madhusudhan et al., 2025) and the speciation cannot therefore be directly compared with that predicted in our models, they provide first-order constraints on the chemical degrees of freedom permitted in sub-Neptune atmospheres.
@@
The lower C/O ratios observed than predicted by our models leads to another discrepancy—the absence of CO2(g) in our simulations. As highlighted for terrestrial planetary atmospheres (Section 3.1), however, any CO(g) that is stable at high temperatures (here, 3000 K) is partitioned between CO2(g), C(cr). and/or CH4(g) at low temperatures (300–400 K, the temperatures in the photosphere recovered from transmission spectra) depending on the prevailing gas chemistry and total pressure. This implies substantial amounts of CO2(g) could form in the upper atmosphere upon cooling. Therefore, future work should concentrate on how gas-condensed phase equilibria act to modify the speciation of sub-Neptune atmospheres throughout the atmospheric column.
A simplified version of this exercise was carried out by Yang & Hu (2024), in which speciation at the pressures of observation in the atmospheres of K2-18b and TOI-270d were extrapolated to 200 bar. In order to match the CH4/CO2 proportions in the upper atmosphere, these authors inferred H2/H2O ratios at 200 bar to be roughly unity for K2-18b and 0.5 for TOI-270d. Assuming these ratios are representative of those at any magma ocean/atmosphere interface, this would correspond to IW in the range 0 to 1. These O2s are similar to, or more oxidized than, those inferred for small, differentiated rocky bodies in the solar system (Wadhwa, 2008). Such high oxygen fugacities would seem at odds with the substantial quantities (1%–3% by mass) of solar-like gas inferred to have been accreted by sub-Neptunes. Indeed, we predict C/H ratios that are too high (2.5) at these O2s, and H2 would no longer be a predominant constituent of the atmosphere (Figure G1). Therefore, preliminary indications suggest that either (1) bulk planetary C/H ratios should be relatively low (10-1 by mass), either by processes that fractionate C from other atmosphere-forming elements during accretion relative to solar or its preferential sequestration into sub-Neptune cores, and/or (2) the lower atmosphere has a distinct chemical composition with higher C/H and C/O than does the upper atmosphere sensed by transmission measurements (e.g., by inefficient mixing or condensation of graphite in the atmospheric column), in order to maintain simultaneously low C/H (0.1–0.3) and C/O (0.02–0.5) ratios observed in temperate sub-Neptunes.